# DistRL: An Asynchronous Distributed Reinforcement Learning Framework for On-Device Control Agents

### v3

@@ 1 Introduction
The integration of Large Language Models (LLMs) into agents capable of complex tasks has gained momentum with initiatives like AutoGPT (Yang et al., 2023a), HuggingGPT (Shen et al., 2024), and MetaGPT (Hong et al., 2023), AutoUI (Zhan & Zhang, 2023), etc. These LLM-based agents extend beyond language processing to perform sophisticated functions, including software development and gaming, leveraging their reasoning abilities to interact with and manipulate environments effectively.
@@
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications (Alayrac et al., 2022, Achiam et al., 2023, Zheng et al., 2024, Li et al., 2023). This versatility enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios (Yang et al., 2023b, Wang et al., 2024). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
However, a gap remains between LLMs’ general reasoning capabilities and their effectiveness in GUI-based device control. While LLMs can process information, they struggle with rational behavior and error recovery in real-world environments. Previous solutions use complex wrappers, but without updating model weights, their performance remains limited (Bai et al., 2022a). Consequently, prior work in building device agents often relies on constructing complex wrappers around these models, combining them with prompting, search, or tool use. However, without updating the model’s weights, the effectiveness of these agents remains inherently limited by the capabilities of the base model.
To bridge this gap, fine-tuning(type of light training) aforementioned agents on demonstrations via imitation learning (Zhan & Zhang, 2023, Nakano et al., 2021, Yang et al., 2023b) channels pre-trained abilities into actionable behaviors suitable for device control. However, the dynamic nature of devices and the web renders models trained on static data sub-optimal as ecosystems evolve. Such agents struggle to recover from mistakes in changing environments, limiting their practical utility. Therefore, building robust and reliable device-control agents necessitates an interactive approach that enables MLLMs to adapt and learn from their own experiences on devices and the Internet.
Moreover, MLLMs are prone to confidently generating incorrect content that deviates from human preferences (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022a, Stiennon et al., 2020). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from Human Feedback (RLHF), becomes essential. RLHF leverages human feedback to align model outputs with desired behaviors and preferences. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to human expectations. Details of the problem formulation as an RL-based fine-tuning approach will be discussed in Sections 2.2 and 3.
One significant challenge in RL-based fine-tuning of mobile agents is the lack of support for efficient online fine-tuning. Existing offline datasets, such as AitW (Rawles et al., 2024b) and AndroidControl (Li et al., 2024a), provide static data that do not capture the dynamic and evolving nature of mobile apps. Frequent updates and new elements like advertisements cause distribution shifts that offline-trained agents struggle to handle, leading to failures in real-world deployments.
@@
The second challenge is the need for Reinforcement Learning (RL) algorithms that can operate efficiently within a distributed framework. Asynchronous data collection introduces algorithmic difficulties: non-stationary data distributions hinder convergence, and delays between policy updates and data collection can cause agents to act on outdated policies, degrading performance. Ensuring consistency and stability becomes more complex when dealing with distributed agents collecting data at different rates and times, necessitating robust mechanisms to handle delayed or out-of-order updates.
To further demonstrate these challenges and the non-trivial nature of bringing MLLMs into mobile device control scenarios, our extensive case studies reveal that even the most advanced proprietary Multimodal Large Language Models (MLLMs) like GPT-4V (Achiam et al., 2023), agents such as AutoUI with Supervised Fine-Tuning (SFT), and state-of-the-art mobile control agents like DigiRL (Bai et al., 2024) fail in numerous scenarios. A detailed analysis of these failure modes is provided in Appendix A.1.
@@
1. Scalable and Asynchronous Architecture for Data Acquisitions: DistRL introduces a decoupled and asynchronous architecture that deploys a tailor-made RL fine-tuned agent across a variety of heterogeneous worker devices and environments for remote data collection. Each worker asynchronously sends real-time interaction data back to the central learner, which continuously updates the agent. This design improves training efficiency and scalability, making DistRL highly effective for adapting control agents to dynamic environments.
2. Advanced RL Algorithm Tailored for Centralized Training: DistRL leverages a novel off-policy reinforcement learning algorithm, A-RIDE (detailed in Section 5), specifically designed for distributed and asynchronous data utlizations. Our algorithm prioritizes significant experiences to enhance sample efficiency, ensuring that the learning process focuses on the most informative data while simultaneously encouraging exploratory behavior among workers.
In practice, we validate our framework using a T5-based multimodal generation model with 1.3B parameters (details in Appendix A.5.1) to efficiently handle both vision and language inputs. To the best of our knowledge, DistRL is the first work to scale autonomous, online RL fine-tuning for mobile device control in a distributed environment.
@@ 2.1 Multi-Modal On-device Control Agents
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT (Nakano et al., 2021) employ fine-tuned models for web browsing, while WebAgent (Gur et al., 2023) generates web code using T5. AppAgent (Yang et al., 2023b) and MobileAgent (Wang et al., 2024) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding (Driess et al., 2023, Reid et al., 2024), but without fine-tuning, they are limited by the base models (Driess et al., 2023).
Some works fine-tune VLMs with demonstration data, e.g., AutoUI, CogAgent (Kapoor et al., 2024, Zhan & Zhang, 2023), but static data-trained models may struggle with real-world variability (Jiang et al., 2023). Others use filtered imitation learning with autonomously collected data (Pan et al., 2024, Lai et al., 2024), While DigiRL (Bai et al., 2024) supports on-device RL fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes. Additionally, DigiRL lacks support for efficient distributed learning algorithm designs, further hindering its scalability and performance in dynamic, parallel settings. To address this, our scalable and asynchronous RL fine-tuning pipeline, DistRL, offers an efficient solution for distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences (Stiennon et al., 2020, Ouyang et al., 2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations (Ouyang et al., 2022, Bai et al., 2022a).
Recent advances in MLLMs (Alayrac et al., 2022, Driess et al., 2023, Reed et al., 2022, Li et al., 2023, Zhang et al., 2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022a, Stiennon et al., 2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative (Bai et al., 2022b, Yu et al., 2024, Lee et al., 2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots (Bai et al., 2024, Lee et al., 2023, Yu et al., 2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems (Reed et al., 2022, Liang et al., 2023, Driess et al., 2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods (Driess et al., 2023, Nakano et al., 2021, Reed et al., 2022, Yao et al., 2023). This approach is effective for large-scale applications like Android device control.
@@ 2.3 Scalable and Distributed RL Framework
Scalable reinforcement learning frameworks like Ray RLlib (Liang et al., 2018) enable distributed training by parallelizing policy learning across CPUs and GPUs using the Ray engine (Liang et al., 2017). Ray RLlib supports various algorithms and efficiently manages neural network training, but it assumes that data collection can be simulated or parallelized within the same infrastructure. This assumption limits its applicability to real-world on-device control tasks involving actual mobile devices, where data collection must occur across distributed, heterogeneous devices with varying task durations and network conditions.
Moreoever, frameworks such as IMPALA (Espeholt et al., 2018) and IMPACT (Luo et al., 2019) are optimized for fast simulations, enabling efficient data collection and policy updates in highly parallelized environments. Applying these frameworks to on-device control introduces non-trivial challenges due to the stochastic nature of real-world interactions and the scalability constraints of mobile devices, including heterogeneous device capabilities, variable task execution times, and unreliable communication.
Our approach builds upon the foundational ideas of IMPALA and IMPACT, effectively extending their concepts to accommodate the unique challenges of on-device control. We introduce a decoupled and asynchronous architecture that supports distributed data collection from heterogeneous, real-world devices by implementing sophisticated communication protocols and optimized data utilization strategies. Furthermore, we develop a novel reinforcement learning algorithm specifically designed to handle the stochasticity and scalability inherent in mobile device environments. This algorithm enhances and maximizes the utilization of collected data while maintaining the exploration behavior of workers, addressing critical aspects that prior work utilizing A3C (Mnih, 2016) have largely overlooked.
@@ 4 System Design
DistRL is an asynchronous distributed reinforcement learning framework for scalable and efficient training of mobile agents. By decoupling trajectory collection from policy learning and doing both in parallel, it leverages distributed working machines for CPU-intense agent-environment interactions and GPU servers for policy training. This separation optimizes efficiency, scalability, and resource utilization by aligning tasks with appropriate hardware.
This decoupled and asynchronous design offers several key advantages: it improves scalability as data collection scales naturally and linearly with more working machines providing the mobile environment, even if there are large performance gaps between them; it optimizes resource utilization by assigning tasks to suitable hardware; and it improves policy quality through richer, more diverse datasets from multiple devices, enhancing robustness and generalization capabilities. The details of our system are presented as follows:
@@
The host learner updates the policy with tailored regularization controls (details in Section 5) to encourage workers to explore a broader range of potential actions during task execution. This approach ensures diversity in the collected data, preventing convergence towards homogeneous behaviors, which is especially crucial in dynamic and complex mobile device environments. Additionally, to maximize the use of the large-scale and diverse data collected, and to avoid excessive learning on redundant or similar data, the learner employs priority-based sampling techniques (details in Section 5). These carefully curated design choices not only ensure training efficiency but also enhance the generalization capability of the policy. After updating the policy, the host learner distributes the latest version to the workers, allowing them to interact with their environments based on the latest learning updates. The training process runs continuously, with new experiences from the workers refining the policy, and the updated policy enhancing the workers’ performance.
@@
On the whole, DistRL employs asynchronous RL to address the challenges of online RL in dynamic, real-world settings. Each thread in workers operates independently, executing tasks and generating learning trajectories at its own pace, which accommodates variability in task durations and system latencies. Data produced by the working threads is queued and processed by the host learner, which updates the global policy based on the collected trajectories. The updated policy is asynchronously distributed back to the workers, allowing for independent and non-blocking policy updates. This approach effectively manages temporal discrepancies between the workers and the host learner, ensuring smooth and effective learning across distributed workers. Further details regarding the communication mechanism between the host and workers are elaborated in Appendix A.3.2.
The asynchronous framework design also ensures scalability. With two 96 vCPU machines, it supports up to 32 emulators operating concurrently, and it can scale almost linearly with worker performance to handle large workloads.
@@ 5 Methodology
In this section, we introduce the core reinforcement learning algorithm employed in DistRL to fine-tune RL agents for device control tasks. The inherent issues of limited on-device resources, asynchronous data generation, and distributed constraints necessitate an efficient and scalable framework.
Reinforcement learning in distributed device control environments encounters significant challenges related to policy stability, convergence, and effective exploration. Stable convergence is essential for reliable agent performance, while robust exploratory behavior is crucial for discovering effective control policies in dynamic and complex settings. On-policy algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Advantage Actor-Critic (A2C) (Mnih, 2016) are limited by their reliance on synchronous data collection and policy updates, leading to sample inefficiency and delayed learning in asynchronous, multi-agent environments.
To overcome these challenges, we introduce A-RIDE, an off-policy reinforcement learning algorithm tailored for distributed environments. A-RIDE stands for Advantage-based Retrace Improved by Distributed Prioritized Experience Replay. It enhances exploration efficiency, maintains policy robustness, and improves training efficiency by promoting robust explorative behavior, ensuring policy stability, and prioritizing informative experiences. These advancements enable DistRL to achieve stable and efficient learning in real-world device control tasks.
@@ 5.1 A-RIDE: The Backbone of DistRL
Our method employs advantage-based estimations to refine policy gradient updates, as an extension of Generalized Advantage Estimation (GAE) (Schulman et al., 2015), effectively balancing exploration and exploitation in the learning process. By introducing a trace decay parameter, A-RIDE manages the bias-variance trade-off in advantage calculations, optimizing the stability and convergence of the policy. A-RIDE incorporates enhancements tailored to distributed, asynchronous environments, ensuring robust policy stability and efficient learning in complex device control tasks.
While GAE has proven highly effective in synchronous environments, its reliance on synchronized data collection makes it less suitable for asynchronous settings, such as distributed control tasks on devices. To handle asynchronous trajectory generation in device control tasks, A-RIDE leverages an enhanced Retrace($\lambda$) method for robust off-policy corrections inspired by (Espeholt et al., 2018, Munos et al., 2016). Unlike traditional on-policy algorithms that require synchronous data collection, A-RIDE is designed for distributed environments with asynchronous data.
The Retrace($\lambda$) update is defined as: $Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\delta_{t}$, where the correction term $\delta_{t}$ is calculated as:
$\delta_{t}=\sum_{k=t}^{H}\gamma^{k-t}\left(\prod_{i=t+1}^{k}c_{i}\right)\left[r_{k}+\gamma Q(s_{k+1},a_{k+1})-Q(s_{k},a_{k})\right].$ (1)
Here, $Q(s_{t},a_{t})$ is the estimated action-value function; $\gamma\in[0,1]$ is the discount factor; $H$ is the time horizon; $c_{i}=\lambda\min\left(1,\rho_{i}\right)$ with $\lambda\in[0,1]$ being the trace decay parameter; $\rho_{i}=\dfrac{\pi(a_{i}|s_{i})}{\mu(a_{i}|s_{i})}$ is the importance sampling ratio between the target policy $\pi$ and the behavior policy $\mu$.
To ensure effective exploration within the action space and prevent the generation of nonsensical or invalid commands, we incorporate entropy regularization into the actor loss function. This addresses the challenge inherent in Vision-Language Models (VLMs) where purely random exploration may lead to semantically incoherent actions (Li et al., 2017). The actor loss is defined as:
@@
$\mathcal{L}=-\mathbb{E}_{\mu}\left[\rho_{t}A(s_{t},a_{t})\log\pi(a_{t}|s_{t})\right]-\beta\mathbb{E}_{\mu}\left[\mathbb{H}(\pi(a_{t}|s_{t}))\right]+\lambda\mathbb{E}_{\mu}\left[\mathcal{P}_{\text{invalid}}(a_{t})\right],$ (2)
@@
where the advantage function $A(s_{t},a_{t})$ represents how much better an action $a_{t}$ is compared to the expected reward at state $s_{t}$, given the policy’s understanding of the environment. It can be expressed as the difference between the action-value function $Q(s_{t},a_{t})$ and the state-value function $V(s_{t})$: $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$. The action-value function $Q(s_{t},a_{t})$ estimates the expected return when taking action $a_{t}$ at state $s_{t}$, while the state-value function $V(s_{t})$ estimates the average expected return from state $s_{t}$ under the policy $\pi$. $\mathbb{H}$ is the entropy term, $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini (Reid et al., 2024)), ensuring that only contextually appropriate actions are penalized. The hyperparameters $\beta$ and $\lambda$ are optimized through empirical studies to balance exploration and policy robustness effectively. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness.

### v4

@@ 1 Introduction
The integration of Large Language Models (LLMs) into agents capable of complex tasks has gained momentum with initiatives like AutoGPT (Yang et al., 2023a), HuggingGPT (Shen et al., 2024), and MetaGPT (Hong et al., 2023), etc. These agents extend beyond language processing to perform sophisticated functions, leveraging their reasoning abilities to interact with and manipulate environments effectively.
@@
One of the key factors driving this trend is the advent of Multimodal Large Language Models (MLLMs), which can process diverse inputs such as text, images, audio, and video, thereby significantly expanding the scope of LLM applications (Alayrac et al., 2022, Achiam et al., 2023, Zheng et al., 2024, Li et al., 2023). This versatility also enables MLLM-based on-device control agents—intelligent systems embedded within mobile devices that manage and operate applications to execute user commands seamlessly—to interact more naturally and efficiently with their surroundings, completing more complex tasks that require a deeper understanding of context and the ability to learn from interactions. For instance, agents designed to operate smartphone applications can interpret screenshots from the operating system, demonstrating flexibility and adaptability that make them valuable tools in a wide range of scenarios (Yang et al., 2023b, Wang et al., 2024a). These agents are essential for tasks such as automating app interactions, managing settings, and enhancing user productivity by providing intuitive control over device functionalities.
@@
Nonetheless, a gap remains between LLMs’ general reasoning capabilities and their effectiveness in GUI-based device control. While LLMs can process information, they struggle with rational behavior and error recovery in real-world environments. Consequently, prior work in building on-device agents often relies on constructing complex wrappers, combining them with prompting, search, or tool use. Without updating the model’s weights, the effectiveness of these agents remains inherently limited by the capabilities of the base model (Bai et al., 2022). Moreover, MLLMs are prone to confidently generating incorrect content that deviates from desired outputs (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022). To address these challenges, introducing reinforcement learning (RL)-based fine-tuning methods, such as Reinforcement Learning from AI Feedback (RLAIF) (Yu et al., 2024, Lee et al., 2023), which evolved from Reinforcement Learning from Human Feedback(RLHF), becomes essential. RLAIF leverages AI-generated feedback to align model outputs with intended behaviors and performance criteria. By incorporating RL-based methods, models can learn to optimize policies that not only perform tasks effectively but also adhere to specific expectations.
@@
One significant challenge in RL-based fine-tuning of mobile agents is the lack of support for efficient online fine-tuning. Existing fine-tuning approaches over mobile agents primarily rely on static offline datasets like AitW (Rawles et al., 2024b) and AndroidControl (Li et al., 2024), which fail to capture the dynamic and evolving nature of mobile applications. Frequent updates and new elements like advertisements cause distribution shifts that offline-trained agents struggle to handle, leading to failures in real-world deployments.
The second challenge is the need for Reinforcement Learning (RL) algorithms that can operate efficiently within a distributed framework. Asynchronous data collection introduces algorithmic difficulties: non-stationary data distributions hinder convergence, and delays between policy updates and data collection can cause agents to act on outdated policies, degrading performance.
@@
1. Scalable and Asynchronous Data Acquisition Architecture: DistRL introduces a decoupled and asynchronous framework that deploys RL agents across heterogeneous worker devices for remote data collection, enhancing training efficiency and scalability (§ 4).
2. Advanced RL Algorithm for Centralized Training: We develop A-RIDE, a novel off-policy reinforcement learning algorithm tailored for distributed and asynchronous data utilization, which prioritizes significant experiences to improve sample efficiency while encouraging exploration (§ 5).
In practice, we validate our framework using a T5-based multimodal generation architecture with 1.3B parameters (details in Appendix A.5.1) to efficiently handle both vision and language inputs. To the best of our knowledge, DistRL is the first deployable and scalable autonomous RL fine-tuning system for online mobile device control tasks in a distributed environment.
@@ 2.1 Multi-Modal On-device Control Agents
Modern approaches use pre-trained models with zero or few-shot prompting and fine-tuning for enhanced capabilities. WebGPT (Nakano et al., 2021) employ fine-tuned models for web browsing, while WebAgent (Gur et al., 2023) generates web code using T5. AppAgent (Yang et al., 2023b) and MobileAgent (Wang et al., 2024a) act as drivers, enabling the LLMs to explore and act on mobile device environments. Training multimodal device control agents poses challenges like pixel-level interactions and variability in device ecosystems. Many rely on proprietary Vision-Language Models (VLMs) and wrappers for GUI visual grounding (Driess et al., 2023, Reid et al., 2024), but without fine-tuning, they are limited by the base models (Driess et al., 2023).
Several works fine-tune Vision-Language Models (VLMs) using demonstration data, such as AutoUI and CogAgent (Kapoor et al., 2024, Zhan & Zhang, 2023); however, models trained on static datasets often struggle with the variability of real-world environments (Jiang et al., 2023). Others employ filtered imitation learning with autonomously collected data (Pan et al., 2024, Lai et al., 2024). While DigiRL (Bai et al., 2024) supports on-device reinforcement learning (RL) fine-tuning, it encounters significant inefficiencies in parallel environments. Specifically, DigiRL’s multi-machine setup relies on a fully synchronous data acquisition process, causing faster workers to idle while waiting for slower ones. This approach is impractical in real-world scenarios where task durations can vary by up to 100 times, ranging from seconds to over ten minutes.
Our extensive case studies reveal significant limitations in these prior works: advanced MLLMs like GPT-4V (Achiam et al., 2023), SFT-based agents like AutoUI (Zhan & Zhang, 2023), and state-of-the-art mobile control agents like DigiRL (Bai et al., 2024) exhibit numerous failure modes (detailed analysis in Appendix A.1). In particular, DigiRL’s lack of efficient distributed learning algorithms severely limits its scalability in dynamic, parallel settings. To address these limitations, we introduce DistRL, a scalable and asynchronous RL fine-tuning pipeline designed for efficient distributed mobile control agent training.
@@ 2.2 Reinforcement Learning for On-device Agent Fine-tuning
Reinforcement Learning from Human Feedback (RLHF) is widely used to fine-tune LLMs to align with human preferences (Stiennon et al., 2020, Ouyang et al., 2022). In device control tasks, similar approaches use imitation learning from human-labeled evaluations, but RLHF is labor-intensive due to the need for human annotations (Ouyang et al., 2022, Bai et al., 2022). Recent advances in MLLMs (Alayrac et al., 2022, Reed et al., 2022, Li et al., 2023) show impressive multimodal capabilities but often produce incorrect outputs that deviate from human preferences (Ouyang et al., 2022, Ziegler et al., 2019, Bai et al., 2022, Stiennon et al., 2020). Reinforcement Learning from AI Feedback (RLAIF), using AI labelers as proxies, offers an alternative (Yu et al., 2024, Lee et al., 2023). For on-device tasks, AI evaluators assess task completion using prompts and screenshots (Bai et al., 2024, Lee et al., 2023, Yu et al., 2024).
Previous RL research focused on single-turn tasks, limiting their effectiveness for multi-step problems (Reed et al., 2022, Liang et al., 2023). To address this, we developed a simplified off-policy multi-turn RL algorithm, A-RIDE, which learns from suboptimal online interactions, reducing complexity and accelerating convergence compared to previous value-based methods (Driess et al., 2023, Yao et al., 2023). This approach is effective for large-scale applications like Android device control.
@@ 2.3 Scalable and Distributed RL Framework
Scalable reinforcement learning frameworks like Ray RLlib (Liang et al., 2018) enable distributed training by parallelizing policy learning across CPUs and GPUs. RLlib supports various algorithms and efficiently manages neural network training, but it assumes that data collection can be simulated or parallelized within the same infrastructure, limiting their applicability to real-world on-device control tasks involving heterogeneous mobile devices with varying task durations and network conditions.
Moreover, IMPALA (Espeholt et al., 2018) and IMPACT (Luo et al., 2019) are influential distributed reinforcement learning algorithms, but we are not able to adopt them directly as baselines for our distributed RL system in mobile device control due to their limitations in addressing the unique challenges of this environment. Specifically, these algorithms inadequately handle the fluctuating online experiences inherent in real-world mobile interactions, lack efficient buffer management for distributed cases, offer limited support for off-policy reinforcement learning models, and do not provide the necessary scalability and system optimizations required for distributed mobile control. Implementing them directly would necessitate substantial modifications to manage communication, schedule worker roles, and handle queues and replay buffers effectively. Therefore, we developed our own approach that builds upon the foundational concepts of IMPALA but extends them to meet the specific requirements of on-device control in mobile environments.
@@ 4 System Design
DistRL is an asynchronous distributed reinforcement learning framework for scalable and efficient training of mobile agents. By decoupling trajectory collection from policy learning and doing both in parallel, it leverages distributed working machines for CPU-intense agent-environment interactions and GPU servers for policy training. This separation optimizes efficiency, scalability, and resource utilization by aligning tasks with appropriate hardware. Moreover, such decoupled and asynchronous design offers several key advantages: it improves scalability as data collection scales with more working machines providing the mobile environment, even if there are large performance gaps between them, optimizes resource utilization by assigning tasks to suitable hardware, and it improves policy quality through richer, more diverse datasets from multiple devices, enhancing robustness and generalization capabilities. The details of our system are presented as follows:
@@
The host learner updates the policy using tailored regularization to promote worker exploration and priority-based sampling to efficiently utilize diverse data. Additionally, to maximize the use of the large-scale and diverse data collected, and to avoid excessive learning on redundant or similar data, the learner employs priority-based sampling techniques. Updated policies are then distributed to workers, creating a continuous cycle of experience collection and policy refinement. Detailed algorithmic design is presented in Section 5.
@@
On the whole, DistRL employs asynchronous RL to address the challenges of online RL in dynamic, real-world settings. Each thread in workers operates independently, executing tasks and generating learning trajectories at its own pace, which accommodates variability in task durations and system latencies. Data produced by the working threads is queued and processed by the host learner, which updates the global policy based on the collected trajectories. The updated policy is asynchronously distributed back to the workers, allowing for independent and non-blocking policy updates. Further details regarding the communication mechanism between the host and workers are elaborated in Appendix A.3.2. The asynchronous framework design also ensures scalability. With two 96 vCPU machines, it supports up to 32 emulators operating concurrently, and it can scale almost linearly with worker performance to handle large workloads.
@@ 5 Methodology
In this section, we introduce A-RIDE, the core reinforcement learning algorithm employed in DistRL to fine-tune agents for device control tasks in distributed environments characterized by limited on-device resources, asynchronous data generation, and distributed constraints. Traditional on-policy algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Advantage Actor-Critic (A2C) (Mnih, 2016) are inefficient in these settings due to their reliance on synchronous data collection and policy updates, leading to sample inefficiency and delayed learning. Unlike existing off-the-shelf RL models that have been widely applied in various domains (Wang & Yoneki, 2024, Wang & Shi, 2021, Wang et al., 2024b, Sun et al., 2022a, Sun & Wang, 2022, Sun et al., 2022b), our approach, A-RIDE (Advantage-based Retrace Improved by Distributed Prioritized Experience Replay), addresses these challenges through two key novel components: Retrace (§ 5.3) and DPER (§ 5.2), enhancing exploration efficiency, maintaining policy robustness, and improving training efficiency through robust exploratory behavior and prioritization of informative experiences. This enables DistRL to achieve stable and efficient learning in real-world device control tasks (see Appendix A.4 for more methodology details).
@@ 5.1 A-RIDE: The Backbone of DistRL
In general, our approach enhances policy gradient updates by extending the Generalized Advantage Estimation (GAE) framework (Schulman et al., 2015) to better suit asynchronous, distributed environments common in device control tasks. Learning from GAE to estimate the advantage function directly, we introduce another stable and smart way to estimate advantages by maintaining two separate networks: one for the trajectory-level value estimation ($V_{\text{traj}}$), which serves as the labeler, and one for the state value function $V(s)$.
Trajectory-Level Value Estimation
We maintain a trajectory-level value estimator $V_{\text{traj}}$, parameterized by a neural network $\theta$. This estimator serves as a labeler, assigning rewards to trajectories and filtering the replay buffer to retain only high-value trajectories. The trajectory-level value estimator is trained using a maximum likelihood estimation (MLE) loss function:
$\min_{\theta}\mathcal{L}(V_{\text{traj}})=-\mathbb{E}_{\nu}\left[r(s_{H},a_{H})\log V_{\text{traj}}(s_{H},a_{H})+\left(1-r(s_{H},a_{H})\right)\log\left(1-V_{\text{traj}}(s_{H},a_{H})\right)\right]$
where $r(s_{H},a_{H})$ is the reward at the time horizon step $H$, and $\nu$ denotes the distribution over trajectories. This estimator helps identify and retain valuable trajectories for more efficient training.
State-Value Function Estimation
We also maintain a separate value network $V(s_{t};\phi)$, which estimates the state value function $V(s_{t})$, representing the expected return from state $s_{t}$. Instead of directly regressing on scalar returns, we train the value network to predict the probability that the Monte Carlo return $G_{t}$ is positive, transforming value estimation into a binary classification task.
The value network is trained with the following loss function: $\mathcal{L}(V)=\mathbb{E}[-\mathbb{I}[G_{t}>0]\log V(s_{t};\phi)-(1-\mathbb{I}[G_{t}>0])\log(1-V(s_{t};\phi))]$, where $V(s_{t};\phi)$ is the predicted probability that $G_{t}>0$, $G_{t}=\sum_{k=t}^{H}\gamma^{k-t}r_{k}$ is the Monte Carlo return from timestep $t$, and $\mathbb{I}[G_{t}>0]$ is an indicator function that equals 1 if $G_{t}>0$ and 0 otherwise. The value network parameters $\phi$ are optimized by: $\phi^{*}=\arg\min_{\phi}\mathcal{L}(V;\phi).$
Advantage Computation
With the trajectory-level rewards and state value estimates obtained, we compute the advantage function $A(s_{t},a_{t})$ using one-step estimation: $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})=r(s_{t},a_{t})+\gamma V(s_{t+1})-V(s_{t})$, which correctly represents the advantage function as per the policy gradient theorem. Here, $r(s_{t},a_{t})$ includes signals of immediate rewards (see § A.4.2). The advantage and value functions are further modified by off-policy corrections, which will be elaborated in Section 5.2.
Policy Optimization with Robust Regularization
Finally, the policy network optimizes the following loss:
@@
$\mathcal{L}=-\mathbb{E}_{\mu}\underbrace{[\rho_{t}A(s_{t},a_{t})\log\pi(a_{t}|s_{t})]}_{\text{Update direction and fine-tuning loss}}-\beta\,\mathbb{E}_{\mu}\underbrace{[\mathbb{H}\left(\pi(a_{t}|s_{t})\right)]}_{\text{Regularization term}}+\lambda\,\mathbb{E}_{\mu}[\underbrace{\mathcal{P}_{\text{invalid}}(a_{t})}_{\text{Action Penalty}}]$ (1)
@@
where $\rho_{t}=\pi(a_{t}|s_{t})/\mu(a_{t}|s_{t})$ is the importance sampling ratio between the target policy $\pi$ and the behavior policy $\mu$, $\mathbb{H}$ is the entropy term for prevention of overfitting (see § 5.3), $\mathcal{P}_{\text{invalid}}(a_{t})$ imposes a penalty on actions deemed invalid based on task-specific criteria, $\beta$ controls the strength of entropy regularization, and $\lambda$ modulates the penalty’s influence. The penalty is assigned using validation through pre-trained LLMs like Gemini (Reid et al., 2024)), ensuring that inappropriate actions are penalized. This formulation encourages the agent to explore a diverse set of actions while constraining it to generate valid and meaningful commands, thereby enhancing both exploration and policy robustness when dealing with online non-stationarities.
This formulation enhances our method over traditional GAE and DigiRL (Bai et al., 2024) by: Incorporating Importance Sampling: The term $\rho_{t}$ ensures that the policy updates remain stable even when learning from off-policy data, which is common in asynchronous, distributed environments. Adding Entropy Regularization: The entropy term $\mathbb{H}(\pi(a_{t}|s_{t}))$ encourages more explorations. Penalizing Invalid Actions: The penalty term $\mathcal{P}_{\text{invalid}}(a_{t})$ discourages the selection of inappropriate or nonsensical actions. Details of achieving the penalties can be found in Appendix A.4.2.
5.2 Off-policy correction: Retrace
To enhance the estimation of the state-value function $V(s_{t})$ in off-policy and asynchronous settings, we apply Retrace($\lambda$) corrections directly to $V(s_{t})$. The Retrace algorithm extends the TD($\lambda$) method to off-policy learning by incorporating importance sampling ratios and a trace decay parameter $\lambda$. Specifically, we update $V(s_{t})$ using the correction term $\delta_{t}$, computed as: $V(s_{t})\leftarrow V(s_{t})+\delta_{t}$, $\delta_{t}=\sum_{k=t}^{H}\gamma^{k-t}(\prod_{i=t+1}^{k}c_{i})[r_{k}+\gamma V(s_{k+1})-V(s_{k})]$, where $c_{i}=\lambda\min\left(1,\rho_{i}\right)$ with $\lambda\in[0,1]$ being the trace decay parameter, and $\rho_{i}$ is the importance sampling ratio as mentioned before. This correction term effectively adjusts the value estimates using future rewards and importance sampling, enabling off-policy learning while mitigating variance due to importance weights. By applying Retrace($\lambda$), we improve the estimation of $V(s_{t})$ in off-policy settings, enhancing the stability and convergence of the value network.